A company produces two types of solar panels per​ year: x thousand of type A and y thousand of type B. The revenue and cost​ equations, in millions of​ dollars, for the year are given as follows.

R(x,y)=3x+5y

​C(x,y)=x^2− 4xy + 9y^2 + 7x - 43y - 9

Determine how many of each type of solar panel should be produced per year to maximize profit.
Answer #1

Reary) = 39+ 54 Clary) = 4y +9y² + 70-434-9 profit PCA, y) = R(x,y)-(Gy) 1. pinig) = fry 72-99 -47 +48449 a = 44-27 -4 ; = 42